OpenStudy (zubhanwc3):

What is the largest possible value for the slopes of all tangent lines to the graph of y=-cos((2x)/3)?

OpenStudy (zubhanwc3):

\[y=-\cos(\frac{ 2x }{ 3 })\]

OpenStudy (zubhanwc3):

my guess, is that when u take the derivative, u get \[y'=\frac{ 2\sin(\frac{ 2x }{ 3 }) }{ 3 }\]and that x should be

OpenStudy (zubhanwc3):

\[\frac{ 3\pi }{ 4 }\] however this is wrong

Parth (parthkohli):

The question doesn't ask you for the value of \(x\) for which the thing is maximum. It asks you what the maximum is, which is 2/3.

OpenStudy (zubhanwc3):

oooo that makes sense, thank you

Parth (parthkohli):

:)